# Towards the Training of Deeper Predictive Coding Neural Networks

### v2

@@ 5 Experiments
To better test the performance of our methods, as well as all the combinations and ablation studies, we report a comprehensive comparison in Table 1. The results show that for shallow networks, we can either match or approximate the performance of BP with all the methods. In deeper models, this is not the case, as we observe that standard PC performs badly, decaying precisions and spiking precisions slightly better, while forward updates well mitigate the degradation of performance. It is however the combination of both inference and update methods that performs the best, with spiking precisions and forward updates always slightly matching the performance of models trained with BP. To conclude, we note that batch freezing further improves the results, clearly showing that the best combination of methods is batch freezing, forward updates, and spiking precisions, which get the best results on all benchmarks when testing on a VGG10. In this case, we use models with normal BN when testing with BP.
@@
To see how much our methods scale on larger datasets and deeper networks, we have also tested them on a VGG15 on the Tiny ImageNet (Le and Yang, 2015), a scaled down version of ImageNet with $200$ classes. The model that we use is identical to the one proposed in the hybrid equilibrium propagation work (Nest and Ernoult, 2024). The results, reported in Table 2, show even more strongly the effectiveness of our method with respect to the original formulation of PC. More in detail, we observe that most of the inference methods alone, even the ones that use precision weighting, fail to reach the performance of $BP$ on this challenging task. However, the results are matched when we use in combination all of the novel methods presented in this work: forward updates, spiking precisions, and BarchNorm freezing.
Algorithm Top-1 Accuracy Top-5 Accuracy
VGG15 VGG15BF VGG15 VGG15BF
BP $\mathbf{44.52^{\pm 0.16}}$ $\mathbf{53.21^{\pm 0.39}}$ $\mathbf{69.28^{\pm 0.06}}$ $\mathbf{77.26^{\pm 0.17}}$
S+F $42.51^{\pm 0.18}$ $53.04^{\pm 0.36}$ $66.22^{\pm 0.18}$ $76.64^{\pm 0.23}$
PC $22.95^{\pm 1.50}$ $22.91^{\pm 0.61}$ $47.04^{\pm 2.04}$ $45.64^{\pm 0.63}$
Decaying Precision (D) $27.29^{\pm 0.24}$ $22.05^{\pm 0.16}$ $54.18^{\pm 0.37}$ $45.22^{\pm 0.23}$
Forward Update (F) $31.61^{\pm 0.73}$ $28.64^{\pm 0.11}$ $55.99^{\pm 0.49}$ $53.72^{\pm 0.31}$
D+F $21.95^{\pm 0.20}$ $30.83^{\pm 0.77}$ $45.15^{\pm 0.23}$ $56.06^{\pm 0.85}$
@@ 6 Conclusion
In this work, we have tackled the problem of scaling up predictive coding for image classification tasks. Specifically, we investigated the following research question: Why do deep models trained with the predictive coding energy fail to match the accuracy of their counterparts trained with backpropagation? We have addressed this by proposing two novel regularization techniques and showed that the combination of our approaches allows the training of deep predictive coding models that come close to the performance of backprop-based models on the same task on deep models. More in detail, we have trained a model with $15$ hidden layers on Tiny Imagenet, a result previously obtained only by using hybrid models that have layers trained with both BP and EP. We believe that these results will inspire future work directed towards making this class of models work at scale, that will tackle more complex datasets and modalities that we have considered, such as the training of equilibrium propagation-based ResNets on ImageNet (He et al., 2016), or small transformer models.
Despite being a good starting point towards the scaling up of the performance of predictive coding and equilibrium propagation models, this work still has some limitations that will be addressed and better studied in future works. The first one relates to the derivation of a better understanding of the phenomenon of the energy imbalance, still not well studied: what causes it, and the correlation it has with the performance of the model on the test and train accuracy. This would allow us to scale to even larger datasets and benchmarks. A second limitation of the work is the presence of the value of the predictions at initialization during the forward updates, which means that the algorithm has to store this value in memory, adding a degree of biological implausibility. This sheds some light on another problem of this class of models, still caused by the energy imbalance: in later layers of the network, the predictions at convergence tend to diverge too much from the predictions at initialization, causing the consequent weight update to be suboptimal. Future work will investigate all of this, and use the derived knowledge to further improve the performance of this class of models, with a nice starting point being a contemporaneous study that theoretically shows it could be possible to train very deep feedforward models Innocenti et al. (2025). Another contemporary work has tackled the same problem as ours by introducing sequential updates for the value nodes Goemaere et al. (2025), similar to a previously introduced variation of PC, that studied the equivalence between PC and backprop Song et al. (2020); Salvatori et al. (2022b).
Acknowledgments
This research was funded in part by the AXA Research Fund and in part by the Austrian Science Fund (FWF) 10.55776/COE12, and in part by VERSES AI. For open access purposes, the author has applied a CC BY public copyright license to any author accepted manuscript version arising from this submission.
@@ Acknowledgments
Appendix A Experiments Setting
@@
We conducted experiments on four models: VGG5, VGG7, VGG10, and VGG15. The detailed architectures of these models are presented in Table 3.
Table 3: Detailed architectures of base models
@@
Channel Sizes [128, 256, 512, 512] [128, 128, 256, 256, 512, 512]
Kernel Sizes [3, 3, 3, 3] [3, 3, 3, 3, 3, 3]
Strides [1, 1, 1, 1] [1, 1, 1, 1, 1, 1]
Paddings [1, 1, 1, 0] [1, 1, 1, 0, 1, 0]
Pool window 2 × 2 2 × 2
Pool stride 2 2
Linear Layers 1 1
@@
Channel Sizes [64, [128]x3, [256]x4, 512] [64, 64, 128, 128, [256]x3, [512]x6]
Kernel Sizes [3, 3, 3, 3, 3, 3, 3, 3, 3] [3, 3, 3, 3, 3, 3, 3, 3, 3, 3, 3, 3]
Strides [1, 1, 1, 1, 1, 1, 1, 1, 1] [1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1]
Paddings [1, 1, 1, 1, 1, 1, 1, 1, 1] [1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1]
Pool window 2 × 2 2 × 2
Pool stride 2 2
Linear Layers 1 2
@@
The benchmark results of above models are obtained with CIFAR10, CIFAR100 and Tiny ImageNet, The datasets are normalized as in Table 4.
Table 4: Data normalization
@@
Table 5: Hyperparameters search configuration
@@
For the optimizer and scheduler, we employ mini-batch stochastic gradient descent (SGD) with momentum for updating $x$ during the relaxation phase. In the Spiking Decaying approach, the learning rate for the last hidden node $x$ decays exponentially with iteration $t$ according to $lr=lr_{x}^{t}$. While we evaluated this exponential decay strategy across other methods, it yielded no performance improvements, leading us to exclude it from their final implementations. For the learning phase, we optimize weights $W$ using AdamW with weight decay. The learning rate schedule follows a warmup-cosine annealing pattern without restarts. This scheduler initiates training with a low learning rate during the warmup period, then smoothly transitions to a cosine-shaped decay curve, preventing abrupt performance degradation. The schedule parameters are configured as follows: the peak learning rate reaches 1.1 times the initial rate, the final learning rate settles at 0.1 times the initial rate, and the warmup phase spans 10% of the total iteration steps.
@@
We conduct hyperparameter selection based on the search space specified in Table 5. All experiments were implemented using the PCX library Pinchetti et al. [2024], a JAX-based framework specifically designed for predictive coding networks that provides comprehensive benchmarking capabilities. All the experiments were conducted on NVIDIA A100/H100 GPUs, with each trial involving a hyperparameter search using the Tree-Structured Parzen Estimator (TPE) algorithm Watanabe [2023] over 500 iterations. The results presented in Table 1 and Figure 4 are obtained using 5 different random seeds (selected from 0-4) with the optimal hyperparameter configuration. The training process is capped at 50 epochs, with an early stopping mechanism that terminates training if no accuracy improvement is observed for 10 consecutive epochs. To maintain consistency with the hyperparameter search settings, we employ a two-phase learning rate schedule: during the first 25 epochs, the weight learning rate follows a warmup-cosine-annealing schedule as previously described, after which it remains fixed at the final learning rate of the scheduler. For the results shown in Figure 2 and  5, we utilize a single random seed with the optimal hyperparameters, setting the maximum training epochs to 50 without implementing early stopping. The weight learning rate schedule remains identical to the aforementioned approach.
(a) VGG-5
(b) VGG-7
Figure 5: Layer-wise Energy Distribution and Accuracy Comparison between BP, PC and Decaying Precision/Spiking Precision with Forward Update in VGG5(a) and VGG7(b) on the CIFAR10 dataset. The colored lines represent the total energy of the individual layers of the model (or, the squared error of every layer for BP). The vertical lines represent the train and test accuracies of the model.
A.1 Computational Complexity
@@
Table 6: Comparison of the training times (seconds) of BP against PCNs on different architectures and datasets.
$\mathbf{BP}$ $\mathbf{PC}$ $\mathbf{S+F}$
VGG-5 - CIFAR10 (T = 5) $1.16^{\pm{0.02}}$ $2.59^{\pm{0.01}}$ $2.37^{\pm{0.01}}$
VGG-7 - CIFAR10 (T = 7) $1.29^{\pm{0.02}}$ $3.75^{\pm{0.01}}$ $3.58^{\pm{0.01}}$
VGG-10 - CIFAR100 (T = 10) $2.03^{\pm{0.06}}$ $10.38^{\pm{0.01}}$ $10.88^{\pm{0.02}}$
VGG-15 - TinyImageNet (T = 15) $9.40^{\pm{0.07}}$ $79.61^{\pm{0.08}}$ $77.58^{\pm{0.02}}$
@@
In Table 6, we present the average time required to train one epoch using $BP$, $PC$, and Spiking Precision + Forward Update ($S+F$) across various tasks on a single H100 GPU. To eliminate the overhead associated with loading datasets into memory, we began timing from the fifth epoch onward, calculating the average duration across five consecutive epochs. We repeated this measurement process five times and report the mean and standard deviation of these five experimental runs. It is worth noting that the reported times for predictive coding suffer from an implementation bottleneck: despite the possibility of updating all the neural activities in parallel, our library does not allow that. This largely slows down our models when trained on deep architectures.
@@ Results.
From the table, we observe that $PC$ and our proposed $S+F$ method have comparable training times, but both are slower than $BP$, with this ratio increasing as the number of model layers increases, mostly for the bottleneck just described. During the forward pass, $BP$ and $PCNs$ ($PC$, $S+F$) perform identical computations, resulting in equivalent computational complexity. However, during the backward pass, $PCNs$ require updating the neuronal activations $x$ before updating weights $W$, with $x$ needing $T$ iterations of updates. The fact that the optimal value of $T$ is always equal to or slightly larger than the number of layers further explains why $PCNs$ are slower compared to $BP$ when the model depth increases, despite this not being as much of a bottleneck as the full parallelization of the operations.
Appendix B Energy Propagation in models
In this section, we deliberately selected the VGG5 and VGG7 architectures from our experimental framework for comparative energy distribution visualization. Figure 5 and Figure 2 present layer-wise energy distribution and classification accuracy among Backpropagation ($BP$), Predictive Coding ($PC$), our proposed Decaying Precision + Forward Update ($D+F$), and our proposed Spiking Precision + Forward Update ($S+F$) evaluated on VGG5 (Figure 5(a)), VGG7 (Figure 5(b)) and VGG10 (Figure 2) architectures using the CIFAR10 dataset.
In VGG5, while all four methods demonstrate comparable accuracy, our proposed methods achieve superior energy equilibrium across network layers compared to $PC$, thereby demonstrating their effectiveness in enhancing the model’s layer-wise energy balance. Interestingly, $PCNs$ achieves good results despite having a less uniform energy distribution than $BP$. Compared to energy propagation in VGG5, Figure 5(b) and Figure 2 reveal a significant energy imbalance in the standard $PC$ when implemented in VGG7/10 architectures, which correlates with degraded test accuracy. In contrast, our proposed method maintains superior energy balance compared to standard $PC$ in deep networks while achieving comparable—or even superior—test accuracy relative to $BP$.
Appendix C Ablation Study

### v3

@@ Results.
We report a comprehensive comparison in Table 2. As expected, in shallow models all methods can either match or approximate the performance of BP with all the methods, while in deeper models, this is not the case for our baselines, but it is for our newly proposed methods. For standard PC, spiking precisions and forward updates alone slightly improve the performance, and it is the combination of both inference and update methods that performs the best, always matching the performance of models trained with BP. To conclude, we note that batch freezing further improves the results, clearly showing that the best combination of methods is batch freezing, forward updates, and spiking precisions, which get the best results on all benchmarks when testing on a VGG10. In this case, we use models with normal BN when testing with BP.
@@
When performing the analysis that lead to the numbers in Table 2, we noted that iPC often performed better without BF. To explicitate this phenomenon, we have performed additional experiments on a VGG15 on the Tiny ImageNet dataset. The model that we use is identical to the one proposed in the hybrid equilibrium propagation work (Nest & Ernoult, 2024). The results, reported in Table 3, show that the effect of BF is much stronger when training models with our version of PC, rather than iPC, whose performance is better when using a vanilla model, without any kind of normalization. Again, the numbers reported here are the result of an hyperparameter search, whose details are in the supplementary material.
Table 3: Test accuracies on Tiny ImageNet.
% Accuracy BP PC+S+F iPC+S
VGG-15
Top-1 $44.52^{\pm 0.16}$ $50.10^{\pm 0.16}$ $\mathbf{50.24^{\pm 0.13}}$
Top-5 $69.28^{\pm 0.06}$ $73.23^{\pm 0.18}$ $\mathbf{75.41^{\pm 0.10}}$
VGG-15-BN/BF
Top-1 $\mathbf{53.21^{\pm 0.39}}$ $53.04^{\pm 0.36}$ $48.19^{\pm 0.71}$
Top-5 $\mathbf{77.26^{\pm 0.17}}$ $76.64^{\pm 0.23}$ $72.64^{\pm 0.20}$
@@ 6 Conclusion
In this work, we have investigated the following research question: Why do deep models trained with the predictive coding energy fail to match the accuracy of their counterparts trained with backpropagation? We have addressed this problem from both the algorithmic and the architectural sides. Algorithmically, we have proposed both a novel technique that leverages a dynamical precision-weighting of prediction errors to better regularize the energy landscape, and a novel weight update mechanism. From the architectural side, we have shown how to modify the residual connections to allow the training of PC-based ResNets, and developed a more effective normalization technique. The results show that we are now able to train VGG models with $15$ layers, and ResNet18 on Tiny Imagenet.
A limitation of the work is the presence of the value of the predictions at initialization during the forward updates, which means that the algorithm has to store this value in memory, adding a degree of biological implausibility, due to computations not being completely local in time anymore. Future work will investigate how to address this problem with a bio-plausible weight update, with a nice starting point being a contemporaneous study that theoretically shows how to train very deep feedforward models (Innocenti et al., 2025). Despite this limitation, however, we have reached the same performance using iPC with spiking updates, and can hence claim that we have reached our goal of training complex models such as ResNet18 with a learning algorithm that is local in time and space, reaching the same performance as backprop.
7 Reproducibility Statement
All the experiments in this work have been run using the PCX library Pinchetti et al. (2024), an open-source software that allows training and testing predictive coding models. Besides this, all the details needed to reproduce the results are carefully described in the supplementary material. We will release the code in case of acceptance.
@@ 7 Reproducibility Statement
LLM Usage.
We have used LLMs for proofreading the paper and to polish writing, for retrieval and discovery of related work, and for low-level coding help (e.g. to help us produce the code for individual figures).
Appendix A Algorithm
The pseudocode is provided in Algorithm 1.
The proposed training procedure is detailed in Algorithm 1. This algorithm presents a unified framework for training both Predictive Coding (PC) and Incremental Predictive Coding (iPC) models, incorporating proposed techniques such as Precision Schedules and Forward Update. The entire process is structured into three core phases:
Initialization: The network’s neuronal activities are initialized through a single feed-forward pass.
Inference Learning (Relaxation Phase): With the input and target values clamped, the hidden layer activities ($x^{l}_{t}$) iteratively relax over T steps to minimize prediction errors across the hierarchy. For iPC, the weights($\mathbf{W}^{l}_{t}$) and neural activities ($x^{l}_{t}$) are jointly updated within each iteration of this relaxation phase.
Weight Update: After the inference phase concludes, the PC model computes and applies its weight updates based on the final states and the initial predictions, following the Forward Update rule.
Algorithm 1 Training a PC/iPC with Precision Schedules and Forward Update
1:Input data $\mathbf{o}$, target $\mathbf{y}$, weights $\{\mathbf{W}^{l}\}_{l=1}^{L}$, activation function $f(\cdot)$
2:Weight learning rate $\eta$, activity learning rate $\alpha$, relaxation steps $T$, precision schedule $\Sigma_{t}^{l}$
3:
4:// Phase 1: Initialization
[…]
6:for $l=1$ to $L$ do $\triangleright$ Initial feed-forward pass
7:  $\bm{\mu}^{l}_{0}\leftarrow\mathbf{W}^{l}f(\mathbf{x}^{l-1}_{0})$
8:  $\mathbf{x}^{l}_{0}\leftarrow\bm{\mu}^{l}_{0}$
9:end for
10:
11:// Phase 2: Inference Learning (Relaxation)
12:Clamp input: $\mathbf{x}^{0}_{t}\leftarrow\mathbf{o}$ for $t\in[1,T]$
13:Clamp output: $\mathbf{x}^{L}_{t}\leftarrow\mathbf{y}$ for $t\in[1,T]$
14:for $t=0$ to T1T-1 do
15:  for $l=1$ to L1L-1 do $\triangleright$ Update hidden layer activities
16:   $\bm{\mu}^{l}_{t+1}\leftarrow\mathbf{W}^{l}f(\mathbf{x}^{l-1}_{t+1})$
17:   $\bm{\varepsilon}^{l}_{t+1}\leftarrow\mathbf{x}^{l}_{t}-\bm{\mu}^{l}_{t+1}$
18:   $\bm{\varepsilon}^{l+1}_{t+1}\leftarrow\mathbf{x}^{l+1}_{t}-\mathbf{W}^{l+1}f(\mathbf{x}^{l}_{t})$
19:   $\Delta\mathbf{x}^{l}_{t}\leftarrow\frac{\alpha}{\Sigma^{l}_{t+1}}\left(\bm{\varepsilon}^{l}_{t+1}-(\mathbf{W}^{l+1})^{\top}\bm{\varepsilon}^{l+1}_{t+1}\odot f^{\prime}(\mathbf{x}^{l}_{t})\right)$
20:   $\mathbf{x}^{l}_{t+1}\leftarrow\mathbf{x}^{l}_{t}+\Delta\mathbf{x}^{l}_{t}$
21:   if iPC then $\triangleright$ Update weights if in iPC Training
22:     $\Delta\mathbf{W}^{l}_{t+1}\leftarrow\frac{\eta}{\Sigma^{l}_{t+1}}\left(\bm{\varepsilon}^{l}_{t+1}f(\mathbf{x}^{l-1}_{t+1})^{\top}\right)$
23:     $\mathbf{W}^{l}_{t+1}\leftarrow\mathbf{W}^{l}_{t+1}-\Delta\mathbf{W}^{l}_{t+1}$
24:   end if
25:  end for
26:end for
27:
28:// Phase 3: Learning (Weight Update for PC with Forward Update)
29:if PC then
30:  for $l=1$ to $L$ do
31:   $\tilde{\bm{\varepsilon}}^{l}_{T}\leftarrow\mathbf{x}^{l}_{T}-\bm{\mu}^{l}_{0}$ $\triangleright$ Calculate Forward Update error
32:   $\Delta\mathbf{W}^{l}\leftarrow\eta\cdot\tilde{\bm{\varepsilon}}^{l}_{T}f(\mathbf{x}^{l-1}_{T})^{\top}$ $\triangleright$ Compute weight update
33:   $\mathbf{W}^{l}\leftarrow\mathbf{W}^{l}-\Delta\mathbf{W}^{l}$ $\triangleright$ Apply weight update
34:  end for
35:end if
Appendix B Experiments Setting
@@
We conducted experiments on four VGG-based models: VGG5, VGG7, VGG10, and VGG15 and two ResNet-based models: ResNet-10, ResNet-18. The detailed architectures of these models are presented in Table 4.
Table 4: Detailed architectures of base models.
@@
Channel Sizes: [128, 256, 512, 512] Channel Sizes: [128, 128, 256, 256, 512, 512]
Kernel Sizes: [3, 3, 3, 3] Kernel Sizes: [3, 3, 3, 3, 3, 3]
Strides: [1, 1, 1, 1] Strides: [1, 1, 1, 1, 1, 1]
Paddings: [1, 1, 1, 0] Paddings: [1, 1, 1, 0, 1, 0]
Pool window: 2 × 2 Pool window: 2 × 2
Pool stride: 2 Pool stride: 2
Linear Layers: 1 Linear Layers: 1
@@
Channel Sizes: [64, [128]x3, [256]x4, 512] Channel Sizes: [64, 64, 128, 128, [256]x3, [512]x6]
Kernel Sizes: [3, 3, 3, 3, 3, 3, 3, 3, 3] Kernel Sizes: [3, 3, 3, 3, 3, 3, 3, 3, 3, 3, 3, 3]
Strides: [1, 1, 1, 1, 1, 1, 1, 1, 1] Strides: [1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1]
Paddings: [1, 1, 1, 1, 1, 1, 1, 1, 1] Paddings: [1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1, 1]
Pool window: 2 × 2 Pool window: 2 × 2
Pool stride: 2 Pool stride: 2
Linear Layers: 1 Linear Layers: 2
ResNet10 ResNet18
Initial Conv: 3x3, Stride 1, Channel Size 64 Initial Conv: 3x3, Stride 1, Channel Size 64
Res-Block: [1, 1, 1, 1] Res-Block: [2, 2, 2, 2]
Channel Sizes: [64, 128, 256, 512] Channel Sizes: [64, 128, 256, 512]
Strides: [1, 2, 2, 2] Strides: [1, 2, 2, 2]
Linear Layers: 1 Linear Layers: 1
@@
The benchmark results of above models are obtained with CIFAR10, CIFAR100 and Tiny ImageNet. The datasets are normalized as in Table 5.
Table 5: Data normalization.
@@
Table 6: Hyperparameters search configuration.
@@
For the optimizer and scheduler, we employ mini-batch stochastic gradient descent (SGD) with momentum for updating $x$ during the relaxation phase. For the learning phase, we optimize weights $W$ using AdamW with weight decay. The learning rate schedule follows a warmup-cosine annealing pattern without restarts. This scheduler initiates training with a low learning rate during the warmup period, then smoothly transitions to a cosine-shaped decay curve, preventing abrupt performance degradation. The schedule parameters are configured as follows: the peak learning rate reaches 1.1 times the initial rate, the final learning rate settles at 0.1 times the initial rate, and the warmup phase spans 10% of the total iteration steps.
@@
We conduct a rigorous hyperparameter search for all models, including the baselines, based on the search space specified in Table 6. All experiments were implemented using the PCX library, a JAX-based framework specifically designed for predictive coding networks that provides comprehensive benchmarking capabilities. All the experiments were conducted on NVIDIA A100/H100 GPUs, with each trial involving a hyperparameter search using the Tree-Structured Parzen Estimator (TPE) algorithm over 200 iterations. The results presented in Table 2 and Figure 3 are obtained using 5 different random seeds (selected from 0-4) with the optimal hyperparameter configuration. The training process is capped at 50 epochs, with an early stopping mechanism that terminates training if no accuracy improvement is observed for 10 consecutive epochs. To maintain consistency with the hyperparameter search settings, we employ a two-phase learning rate schedule: during the first 25 epochs, the weight learning rate follows a warmup-cosine-annealing schedule as previously described, after which it remains fixed at the final learning rate of the scheduler. For the results shown in Figure 25 and 6, we utilize a single random seed with the optimal hyperparameters, setting the maximum training epochs to 50 without implementing early stopping. The weight learning rate schedule remains identical to the aforementioned approach.
Appendix C Computational Complexity
@@
Table 7: Comparison of the training times (seconds per epoch) of BP against PCNs on different architectures with CIFAR10
Task $\mathbf{BP}$ $\mathbf{PC}$ $\mathbf{PC+S+F}$ $\mathbf{iPC}$ $\mathbf{iPC+S}$
VGG5 (T = 5) $1.16^{\pm{0.02}}$ $1.56^{\pm{0.01}}$ $1.55^{\pm{0.01}}$ $1.94^{\pm{0.01}}$ $1.94^{\pm{0.01}}$
VGG7 (T = 7) $1.29^{\pm{0.02}}$ $20^{\pm{0.01}} 2.2$$20^{\pm{0.01}}$ $2.94^{\pm{0.01}}$ $2.95^{\pm{0.01}}$
VGG10 (T = 10) $1.90^{\pm{0.01}}$ $5.08^{\pm{0.02}}$ $5.06^{\pm{0.02}}$ $7.00^{\pm{0.05}}$ $7.01^{\pm{0.03}}$
ResNet10 (T = 10) $1.75^{\pm{0.01}}$ $5.02^{\pm{0.02}}$ $5.10^{\pm{0.01}}$ $7.08^{\pm{0.05}}$ $7.08^{\pm{0.05}}$
ResNet18 (T = 18) $2.78^{\pm{0.01}}$ $14.92^{\pm 0.01}$} $15.17^{\pm{0.03}}$ $2.14^{\pm{0.05}} 22$$2.14^{\pm{0.01}}$
@@
In Table 7, we present the average time required to train one epoch using BP, PC, iPC, iPC with Spiking Precision (iPC + S) and PC with Spiking Precision and Forward Update (PC + S + F) across various tasks on a single H100 GPU. To eliminate the overhead associated with loading datasets into memory, we began timing from the fifth epoch onward, calculating the average duration across five consecutive epochs. We repeated this measurement process five times and report the mean and standard deviation of these five experimental runs. It is worth noting that the reported times for predictive coding suffer from an implementation bottleneck: despite the possibility of updating all the neural activities in parallel, our library does not allow that. This largely slows down our models when trained on deep architectures.
@@ Results.
The results in Table 7 lead to several key observations. First, our proposed methods, PC+S+F and iPC+S, exhibit training times nearly identical to their respective baselines, PC and iPC. This demonstrates that the proposed spiking precision and forward update mechanisms do not introduce a significant computational overhead.
The table also shows that iPC is consistently slower than PC. This performance difference stems from their distinct weight update strategies. PC first runs the inference learning for T timesteps to allow the neural activities ($x$) to converge and then performs a single weight update at the end. In contrast, iPC performs weight updates within each of the T inference steps, alongside the neural activity updates. This approach results in T separate weight updates for every layers instead of one, is the direct cause of its higher computational cost compared to PC.
Finally, all PC models are slower than BP, with this ratio increasing as the number of model layers increases, mostly for the bottleneck just described. While the forward pass is computationally identical for both BP and PC, their backward/update passes differ fundamentally. BP computes gradients in a single backward pass. In contrast, PC perform an iterative inference process to update neural activities by minimizing a global prediction error. This process runs for a fixed number of T. The computational complexity of a single BP epoch is proportional to the number of layers, L. For PC, each of the T inference steps involves computations across all layers, making their complexity roughly proportional to T×L. Since optimal performance often requires T to be equal to or greater than the network depth L, the computational cost for PC naturally scales more rapidly with deeper architectures, despite this not being as much of a bottleneck as the full parallelization of the operations. Thus, We expect predictive coding networks to maintain computational efficiency across larger model architectures while offering substantial performance advantages when implemented on specialized analog neuromorphic hardware.
Appendix D Ablation Study
In this section, we conduct ablation studies to evaluate the individual and synergistic effects of our proposed components. By systematically isolating each mechanism, we quantify its specific contribution to the overall performance of the model.
D.1 Spiking Precision
D.1.1 Comparison of different Precision Schedules
Our central hypothesis is that mitigating the energy imbalance in deep networks requires a potent and precisely timed signal amplification. This amplification should occur at the moment the error information, propagating from the output layer, first arrives at a given hidden layer. To test this hypothesis, we designed and compared several dynamic precision schedules.
Based on this perspective, in addition to Spiking Precision, we designed a Decaying Precision schedule, which offers a slightly smoother, yet still powerful, amplification profile. The formula for Decaying Precision is as follows:
Here, the numerator sum serves as a normalization term that ensures that the sum of the layer-wise precisions over time is equal to one, that is, $\sum_{t=1}^{T}(\Sigma_{t}^{l})^{-1}=1$. The denominator $e^{-k\cdot(l-L+t)}$ allows lower layers to receive larger weights when activated ($l\geq L-t$), thereby helping to achieve a more balanced energy distribution during the inference phase, k is a hyperparameter that controls the strength of this balancing effect, the search range is $[1.0,1.5,2.0]$. It also ensures that each layer experiences a significant boost in precision precisely when the energy from the output first reaches that layer (l=Lt)(l=L-t). When l<Ltl<L-t, we set $\Sigma^{l}_{t}=1$.
Table 8: Test accuracies of the different algorithms across architectures and datasets.
Dataset Algorithm VGG5 VGG7 VGG10 VGG5BF VGG7BF VGG10BF
Decaying Precision (D) $87.91^{\pm 0.22}$ $81.14^{\pm 0.19}$ $4.87^{\pm 0.19} 88$$8.55^{\pm 0.09}$ $80.00^{\pm 0.10}$ $81.68^{\pm 0.18}$
PC+D+F $89.32^{\pm 0.14}$ $89.34^{\pm 0.09}$ $89.43^{\pm 0.18}$ $\mathbf{90.37^{\pm 0.13}}$ $\mathbf{91.48^{\pm 0.12}}$ $91.46^{\pm 0.12}$
PC+S+F $\mathbf{89.45^{\pm 0.18}}$ $\mathbf{90.08^{\pm 0.21}}$ $92.07^{\pm 0.10}$ $89.30^{\pm 0.13}$ $90.89^{\pm 0.04}$ $\mathbf{93.27^{\pm 0.10}}$
CIFAR100 (Top-1) BP $66.28^{\pm 0.23}$ $65.36^{\pm 0.15}$ $\mathbf{69.35^{\pm 0.16}}$ $39^{\pm 0.25} 67.67$$67^{\pm 0.11}$ $71.25^{\pm 0.21}$
PC+D+F $66.10^{\pm 0.09}$ $4.86^{\pm 0.10} 66$$6.54^{\pm 0.12}$ $\mathbf{67.56^{\pm 0.25}}$ $67.27^{\pm 0.21}$ $69.81^{\pm 0.22}$
PC+S+F $\mathbf{66.49^{\pm 0.15}}$ $\mathbf{66.34^{\pm 0.22}}$ $69.08^{\pm 0.08}$ $67.16^{\pm 0.16}$ $\mathbf{67.71^{\pm 0.10}}$ $\mathbf{72.02^{\pm 0.12}}$
[…]
Decaying Precision (D) $81.59^{\pm 0.13}$ $74.00^{\pm 0.30}$ $3.13^{\pm 0.74} 88$$8.82^{\pm 0.07}$ $78.93^{\pm 0.29}$ $81.16^{\pm 0.36}$
PC+D+F $85.85^{\pm 0.10}$ $83.80^{\pm 0.20}$ $86.10^{\pm 0.21}$ $\mathbf{89.84^{\pm 0.17}}$ $89.74^{\pm 0.12}$ $91.24^{\pm 0.07}$
PC+S+F $\mathbf{86.36^{\pm 0.11}}$ $\mathbf{84.53^{\pm 0.15}}$ $86.84^{\pm 0.07}$ $89.57^{\pm 0.09}$ $89.62^{\pm 0.18}$ $\mathbf{92.10^{\pm 0.10}}$
As shown in Table 8, both Decaying and Spiking Precision schedules offer improvements over the baseline PC, particularly in deeper models like VGG10. However, a clear pattern emerges when we analyze their effectiveness in relation to network depth. In shallower models like VGG5 and VGG7, the performance of Decaying Precision is comparable to that of Spiking Precision. This suggests that when the signal path is short, a moderately amplification is sufficient.
Algorithm Top-1 Accuracy Top-5 Accuracy
VGG15 VGG15BF VGG15 VGG15BF
PC+S+Forward Update $42.51^{\pm 0.18}$ $53.04^{\pm 0.36}$ $66.22^{\pm 0.18}$ $76.64^{\pm 0.23}$
PC $2.95^{\pm 1.50} 22$$2.91^{\pm 0.61}$ $47.04^{\pm 2.04}$ $45.64^{\pm 0.63}$
Decaying Precision (D) $7.29^{\pm 0.24} 22$$2.05^{\pm 0.16}$ $54.18^{\pm 0.37}$ $45.22^{\pm 0.23}$
PC+D+Forward Update $21.95^{\pm 0.20}$ $30.83^{\pm 0.77}$ $45.15^{\pm 0.23}$ $56.06^{\pm 0.85}$
As shown in the Tab. 9, when model depth increases to VGG15 with TinyImageNet task, a noticeable performance gap appears, with Spiking Precision consistently outperforming Decaying Precision. This finding strongly supports our core hypothesis: the exponential signal attenuation in deeper networks necessitates a correspondingly sharp and powerful counteracting signal. The abrupt, targeted amplification of Spiking Precision is more effective at preserving the integrity of the error signal across many layers than the smoother profile of Decaying Precision. Consequently, for all subsequent experiments reported in the main body of this paper, we exclusively utilized the superior Spiking Precision schedule. This investigation also opens exciting avenues for future work, such as exploring hybrid schedules that might combine the strengths of different amplification profiles.
D.1.2 Energy Propagation with Precision
We observed that removing the decaying/spiking precision module consistently leads to performance degradation. This effect is particularly evident in deeper models like VGG7 and VGG10, where its absence causes a significant imbalance in the energy distribution across layers. For instance, as shown in Figure 5, the energy proportion of the first layer in the VGG7 model with spiking precision is approximately $10^{-6}$. Without this precision term, the proportion plummets to $10^{-18}$. A comparison of the layer-wise energy distributions (Figure 6) confirms that our proposed precision methods effectively rebalance energy propagation. By increasing the energy in the initial layers by several orders of magnitude, these methods rectify the imbalance, which contributes directly to improved model performance.
Furthermore, the degree of this energy rebalancing correlates with the performance difference between the Decaying and Spiking Precision variants. In the VGG5 and VGG7 models, where the accuracy gap between the two methods is minimal, the difference in their first-layer energy distributions is also small. However, in VGG10, where Spiking Precision significantly outperforms Decaying Precision, the energy gap is far more pronounced. Specifically, the first layer’s energy proportion is approximately $10^{-10}$ for Decaying Precision, whereas Spiking Precision elevates it to $10^{-4}$, highlighting a clear link between balanced energy propagation and model accuracy.